# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ IV-E Decentralized Update of $\mathbf{T}$ and $\underline{\mathbf{R}}$
$[\mathbf{D}_{k}^{\mathtt{Tx}}]^{c}=$ $\sqrt{\alpha_{k}}\overline{\mathbf{W}}_{k}^{c}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}$
@@
$-\sum_{j=1}^{K}\overline{\mathbf{W}}_{j}^{c}\tilde{\mathbf{G}}_{kj}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}.$ (74)
@@
Meanwhile, as indicated by (III-E), calculating $\mathbf{D}_{k}^{\mathtt{Rx}}$ requires only $\tilde{\mathbf{G}}_{kj}$, which has been computed using (48). The remaining calculations in (III-E) are performed at the CU, given by
@@
$-\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\sum_{j=1}^{K}\tilde{\mathbf{G}}_{kj}\tilde{\mathbf{G}}_{kj}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}.$ (75)
Scalars $\delta^{\mathtt{Tx}}$ and $\delta_{k}^{\mathtt{Rx}}$ are stored at the CU. To compute the values distributively, we define $\tilde{\mathbf{W}}_{kj}\triangleq\overline{\mathbf{W}}_{k}^{\mathsf{H}}\overline{\mathbf{W}}_{j}$ and compute $\tilde{\mathbf{W}}_{kj}$ distributively. First, each DU computes $(\overline{\mathbf{W}}_{k}^{c})^{\mathsf{H}}\overline{\mathbf{W}}_{j}^{c}$ and transmits the result to the CU, where $\tilde{\mathbf{W}}_{kj}$ is aggregated as
$\tilde{\mathbf{W}}_{kj}=\sum\nolimits_{c=1}^{C}\left(\overline{\mathbf{W}}_{k}^{c}\right)^{\mathsf{H}}\overline{\mathbf{W}}_{j}^{c}.$ (76)
However, $\delta^{\mathtt{Tx}}$ still cannot be computed distributively since $\hat{\mathbf{W}}$ in (31) cannot be stored and calculated. Therefore, we apply the triangle inequality and Cauchy-Schwarz inequality to $\sum_{j=1}^{M}\big{\lvert}[\hat{\mathbf{W}}]_{mj}\big{\rvert}$ and obtain
@@
To calculate the r.h.s. of (77), each DU first computes $\sum_{j=1}^{M_{c}}\big{\lVert}[\overline{\mathbf{W}}_{k}^{c}]_{j}\big{\rVert}_{2}$ and transmits it to the CU. Then, the CU calculates the summation $\sum_{c=1}^{C}\sum_{j=1}^{M_{c}}\big{\lVert}[\overline{\mathbf{W}}_{k}^{c}]_{j}\big{\rVert}_{2}$ and broadcasts it back to DUs, where the r.h.s. of (77) is finally computed.
Plugging (76) and (77) into (31), we obtain the expression for computing $\delta^{\mathtt{Tx}}$, given by (78) at the bottom of the previous page. To compute $\delta^{\mathtt{Tx}}$ distributively as in (78), the CU first broadcasts $\tilde{\mathbf{W}}_{kj}$ to all DUs. Then, each DU computes the argument of the maximization operation in (31) and selects the greatest value to send to the CU. After receiving $C$ values from the DUs, the CU determines the largest value, and set it as $\delta^{\mathtt{Tx}}$. The expression for $\delta^{\mathtt{Rx}}_{k}$ is obtain by plugging (76) into (43), and it is given by (79) at the bottom of the previous page. The value of $\delta^{\mathtt{Rx}}_{k}$ is computed directly at the CU.
Once the values of $\mathbf{D}_{k}^{\mathtt{Tx}}$, $\mathbf{D}_{k}^{\mathtt{Rx}}$, $\delta^{\mathtt{Tx}}$, and $\delta_{k}^{\mathtt{Rx}}$ are obtained, the surrogate functions in (27) and (38) can be computed accordingly. To update $\mathbf{T}$, the CU first broadcasts $\delta^{\mathtt{Tx}}$ to all DUs. Each DU then computes the closed-form unconstrained global optimum $[\mathbf{T}^{\star}]^{c}$ as
$[\mathbf{T}^{\star}]^{c}=\overline{\mathbf{T}}^{c}+\frac{1}{\delta^{\mathtt{Tx}}}\nabla_{\mathbf{T}^{c}}f_{\rm Quad}\left(\overline{\mathbf{T}}\right),$ (80)
where the entries of $\nabla_{\mathbf{T}^{c}}f_{\rm Quad}(\overline{\mathbf{T}})$ are computed at the DU using (28), with $\mathbf{D}_{k}^{\mathtt{Tx}}$ replaced by $[\mathbf{D}_{k}^{\mathtt{Tx}}]^{c}$. Once $[\mathbf{T}^{\star}]^{c}$ is obtained, the DU updates $\mathbf{T}^{c}$ by projecting $[\mathbf{T}^{\star}]^{c}$ onto the cuboid region using (47a). Moreover, the CU updates $\underline{\mathbf{R}}$ as described in (44) and (47b). To summarize, the key processes of the above decentralized implementation of Algorithm 1 are presented in Algorithm 2.
@@
1:$C$, $M$, $N$, $K$, $P_{\max}$, $\alpha_{k}$, $\mathbf{\Sigma}_{k}$, $L_{k}^{t}$, $L_{k}^{r}$, $\theta_{ki}^{t}$, $\phi_{ki}^{t}$, $\theta_{kj}^{r}$, $\phi_{kj}^{r}$.
@@
5:     steps in Section IV-C and store the results at the CU.
6:     Update each   $[\mathbf{W}_{k}]^{c}$   distributively   according   to   the
7:     steps in Section IV-D and store it at the $c$-th DU.
8:     Update  $[\mathbf{T}]^{c}$  and   $\mathbf{R}_{k}$   distributively according to the
9:     steps in Section IV-E,   and store them at the $c$-th DU
@@
12:$[\mathbf{W}_{k}]^{c}$, $[\mathbf{T}]^{c}$, $\mathbf{R}_{k}$.
IV-F Complexity Analysis
For the centralized implementation given in Algorithm 1, the time complexity of a single BCA iteration is dominated by the updates of $\underline{\mathbf{W}}$ and $\mathbf{T}$, and is given by $\mathcal{O}(M^{3}T_{\rm bis}+M^{2}L_{k}^{t}T_{\rm MM}^{\mathtt{Tx}})$. Here, $T_{\rm bis}$ denotes the number of iterations in the bisection search for updating $\underline{\mathbf{W}}$ described in Section III-C, and $T_{\rm MM}^{\mathtt{Tx}}$ represents the average number of MM iterations required for updating $\mathbf{T}$. The complexity of the centralized implementation is proportional to $M^{3}$ and grows rapidly as the number of transmit FAs $M$ increases.
In comparison, the complexity of the decentralized implementation at the CU is given by $\mathcal{O}(K^{2}d^{3}+N^{3}K+T_{\rm MM}^{\mathtt{Rx}}(N^{2}L_{k}^{r}+N(L_{k}^{r})^{2}))$, and the complexity at each DU is given by $\mathcal{O}(NK^{2}d+K^{2}d^{2}+T_{\rm MM}^{\mathtt{Tx}}(M_{c}dL_{k}^{t}+d(L_{k}^{t})^{2}))$, where $T_{\rm MM}^{\mathtt{Tx}}$ and $T_{\rm MM}^{\mathtt{Rx}}$ represent the MM iterations for the update step of $\mathbf{T}$ and $\underline{\mathbf{R}}$, respectively. Notably, the complexities of the CU and each DU are independent of $M$ and are significantly lower than that of the centralized implementation.
V Simulation
In this section, we evaluate the performance of FA-assisted MU-MIMO networks optimized using the proposed centralized BCA-based algorithm in Algorithm 1 and its decentralized implementation in Algorithm 2, denoted as “C” and “D”, respectively.
We denote the system with joint beamforming and transmit and receive FA position optimization as transmit and receive FA (TRFA), which is optimized by our algorithms. We compare the performance of TRFA with several baselines, specified as follows.
@@ V Simulation
    FPA: The antenna arrays at the BS and users are fixed in position with a spacing of $\lambda/2$. The update steps for $\mathbf{T}$ and $\underline{\mathbf{R}}$ are skipped.
@@
    Transmit FA (TFA): The antennas at the users are fixed with a spacing of $\lambda/2$, while the antennas at the BS are FAs. The update step for $\underline{\mathbf{R}}$ is skipped.
@@
    Receive FA (RFA): The antennas at the BS are fixed with a spacing of $\lambda/2$, while the antennas at the users are FAs. The update step for $\mathbf{T}$ is skipped.
V-A Simulation Setup
We consider a downlink MU-MIMO system with $K=6$ users, each equipped with $N=4$ FAs. The number of parallel data streams is set to $d=4$. All users have equal priority, given by $\alpha_{k}=1$. The system operates over a millimeter-wave channel with a carrier frequency of $f_{c}=28$ GHz [37]. Unless otherwise specified, the power budget is set to $P_{\max}=20$ dBm. The users are uniformly distributed within an annular region centered at the BS. Consequently, the distance between the BS and each UE $d_{k}$ follows the distribution $d_{k}^{2}\sim\mathcal{U}\left[d_{\min}^{2},d_{\max}^{2}\right]$, where $d_{\min}=20$ m and $d_{\max}=100$ m. The pathloss is modeled as $\kappa(d_{k})=T_{0}\big{(}\frac{d_{k}}{d_{0}}\big{)}^{-\varrho}$, where the reference pathloss at $d_{0}=1$ m is $T_{0}=20\lg(4\pi d_{0}/\lambda)=61.4$ dB, and the pathloss exponent is $\varrho=3.67$ [30, 31]. The elevation and azimuth AoAs and AoDs are assumed to follow uniform distributions, i.e., $\theta_{ki}^{t}\sim\mathcal{U}\left[0,\pi\right]$, $\phi_{ki}^{t}\sim\mathcal{U}\left[0,\pi\right]$, $\theta_{kj}^{r}\sim\mathcal{U}\left[0,\pi\right]$, and $\phi_{kj}^{r}\sim\mathcal{U}\left[0,\pi\right]$. The number of transmit and receive paths between the BS and user $k$ is assumed identical, i.e., $L_{k}^{t}=L_{k}^{r}=3$ [37]. Consequently, the PRM matrices $\mathbf{\Sigma}_{k}$ are square and assumed diagonal, with elements distributed as $[\mathbf{\Sigma}_{k}]_{qq}\sim\mathcal{CN}\left(0,(\kappa(d_{k})L)^{-1}\right)$ [25]. Unless otherwise specified, the simulation results are averaged over $200$ independent channel realizations.
The minimal distance between any pair of FAs is set as $D=\lambda/2$. Let $\rho$ denote the movable region of each antenna, normalized by $\lambda$. Without the box-constrained movement mode, all transmit (and receive) FAs can move within a shared cuboid region, where the edge length is $\rho\lambda\sqrt{M}$ (and $\rho\lambda\sqrt{N}$), and the height is $2\rho\lambda$. With the box-constrained movement mode, the specific movable regions of each FA are detailed in Section III-F.
V-B Impact of Box-Constrained Movement Mode

### v2

@@ V-E Decentralized Update of $\mathbf{T}$ and $\underline{\mathbf{R}}$
$\delta^{\mathtt{Tx}}_{m,c}=\frac{24\pi^{2}}{\lambda^{2}}$ $\sum_{k=1}^{K}L_{k}^{\mathtt{Tx}}\Big[\Big(\sum_{t=1}^{K}\Big\lVert[\overline{\mathbf{W}}_{t}^{c}]_{m}\Big\rVert_{2}\sum_{j=1}^{M}\Big\lVert[\overline{\mathbf{W}}_{t}]_{j}\Big\rVert_{2}+\sum_{t=1}^{K}\sum_{s=1}^{K}[\overline{\mathbf{W}}_{t}^{c}]_{m}\tilde{\mathbf{W}}_{ts}[\overline{\mathbf{W}}_{s}^{c}]_{m}^{\mathsf{H}}\Big)\lVert\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Tx}}\rVert_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{\mathtt{Tx}}}}\Big\lVert[\overline{\mathbf{W}}_{k}^{c}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big\rVert_{2}\Big].$ (75)
@@
$\delta^{\mathtt{Rx}}_{k}=\underset{1\leq n\leq N}{\max}\frac{24\pi^{2}}{\lambda^{2}}$ $L_{k}^{\mathtt{Rx}}\left(\left(\sum\nolimits_{j=1}^{N}\Big\lvert[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)[\overline{\mathbf{\Phi}}_{k}]_{j}^{\mathsf{H}}\Big\rvert+\sqrt{N}\Big\lVert[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\Big\rVert_{2}\right)\Big\lVert\sum\nolimits_{t=1}^{K}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kt}\tilde{\mathbf{G}}_{kt}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big\rVert_{2}\right$. (76)
$\left.+\sqrt{\frac{\alpha_{k}}{L_{k}^{\mathtt{Rx}}}}\Big\lVert[\mathbf{\Phi}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big\rVert_{2}\right).$
To update $\mathbf{T}$ and $\underline{\mathbf{R}}$ distributively, we must compute $\mathbf{D}_{k}^{\mathtt{Tx}}$, $\mathbf{D}_{k}^{\mathtt{Rx}}$, $\delta^{\mathtt{Tx}}$, and $\delta^{\mathtt{Rx}}_{k}$ distributively in advance. The matrix $\mathbf{D}_{k}^{\mathtt{Tx}}$ is stored distributively as $\mathbf{D}_{k,c}^{\mathtt{Tx}}$, whereas $\mathbf{D}_{k}^{\mathtt{Rx}}$ is stored at the CU. To calculate $\mathbf{D}_{k}^{\mathtt{Tx}}$, the CU broadcasts $\overline{\mathbf{\Gamma}}_{k}$, $\overline{\mathbf{\Phi}}_{k}$, $\overline{\mathbf{F}}_{k}$, and $\tilde{\mathbf{G}}_{kj}$ to each DU. Then, the $c$-th DU computes $\mathbf{D}_{k,c}^{\mathtt{Tx}}$ as
$\mathbf{D}_{k,c}^{\mathtt{Tx}}=$ $\sqrt{\alpha_{k}}\overline{\mathbf{W}}_{k}^{c}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}$
$-\sum\nolimits_{j=1}^{K}\overline{\mathbf{W}}_{j}^{c}\tilde{\mathbf{G}}_{kj}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}.$ (72)
@@
Meanwhile, as indicated by (IV-E), calculating $\mathbf{D}_{k}^{\mathtt{Rx}}$ requires only $\tilde{\mathbf{G}}_{kj}$. The remaining calculations in (IV-E) are performed at the CU, given by
@@
$-\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\sum\nolimits_{j=1}^{K}\tilde{\mathbf{G}}_{kj}\tilde{\mathbf{G}}_{kj}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}.$ (73)
Scalars $\delta^{\mathtt{Tx}}$ and $\delta_{k}^{\mathtt{Rx}}$ are stored at the CU. To compute the values distributively, we define $\tilde{\mathbf{W}}_{kj}\triangleq\mathsf{Mul}(\overline{\mathbf{W}}_{k}^{c},\overline{\mathbf{W}}_{j}^{c})$ and it is calculated as shown in Fig. 3. However, $\delta^{\mathtt{Tx}}$ still cannot be computed distributively since $\hat{\mathbf{W}}$ in (IV-D) cannot be stored and calculated. We thus apply the triangle inequality and Cauchy-Schwarz inequality to $\sum_{j=1}^{M}\big\lvert[\hat{\mathbf{W}}]_{mj}\big\rvert$ and obtain
@@
To calculate the r.h.s. of (74), each DU first computes $\sum_{j=1}^{M_{c}}\big\lVert[\overline{\mathbf{W}}_{k}^{c}]_{j}\big\rVert_{2}$ and transmits it to the CU. Then, the CU calculates the summation $\sum_{c=1}^{C}\sum_{j=1}^{M_{c}}\big\lVert[\overline{\mathbf{W}}_{k}^{c}]_{j}\big\rVert_{2}$ and broadcasts it back to DUs, where the r.h.s. of (74) is finally computed. Plugging (74) into (IV-D), we obtain the expression for $\delta^{\mathtt{Tx}}$. To better demonstrate the decentralized computation of $\delta^{\mathtt{Tx}}$, we define $\delta^{\mathtt{Tx}}=\underset{1\leq c\leq C}{\max}\underset{1\leq m\leq M_{c}}{\max}\delta^{\mathtt{Tx}}_{m,c}$, and the expression of $\delta^{\mathtt{Tx}}_{m,c}$ is given by (75) at the top of the next page. With $\tilde{G}_{kj}$ computed previously, the value of $\delta^{\mathtt{Rx}}_{k}$ can be directly calculated at the CU, as shown in (76) at the top of the next page.
Here, we summarize the update of $\underline{\mathbf{T}}$ under the DBP architecture. We refer directly to the centralized equations introduced earlier to avoid presenting similar formulas. The only modification in the decentralized setting is to append a superscript $c$ to the variable $\mathbf{t}_{m}$ and a subscript $c$ to $\mathbf{D}_{k}^{\mathtt{Tx}}$. First, we compute the coefficients of the surrogate function $h^{\mathtt{Tx}}(\mathbf{T}\big|\underline{\overline{\mathbf{T}}})$. This begins with the distributed computation of $\tilde{\mathbf{W}}_{kj}$ as illustrated by Fig. 3. After the CU broadcasts $\tilde{\mathbf{W}}_{kj}$ to all DUs, the entries of $\nabla_{\mathrm{vec}(\mathbf{T}^{c})}f_{\rm Quad}(\overline{\mathbf{T}}^{c})$ are calculated by (IV-D), (29), and (V-E) at each DU. Next, to compute $\delta^{\mathtt{Tx}}$, each DU evaluates $\delta^{\mathtt{Tx}}_{m,c}$ by (75) and selects the greatest value to send to the CU, which determines the largest and sets it as $\delta^{\mathtt{Tx}}$. With these coefficients computed, the optimal solution for $\underline{\mathbf{T}}$ is computed at the DUs using (34) and (47a). The aforementioned process is summarized in Fig. 5, where only the $c$-th DU is shown for brevity.
Figure 5: The decentralized update of $\mathbf{T}$ under the DBP architecture.
The calculation of $\mathbf{R}_{kn}$ can be directly executed at the CU. First, we compute the coefficients of the surrogate function $h^{\mathtt{Rx}}_{k}(\mathbf{R}_{k}\big|\overline{\mathbf{R}}_{k})$. Leveraging the $\tilde{\mathbf{G}}_{kj}$ computed previously, the CU computes the entries of $\nabla_{\mathrm{vec}(\mathbf{R}_{k})}f_{\rm Quad}(\overline{\mathbf{R}}_{k})$ and $\delta_{k}^{\mathtt{Rx}}$ using (40) and (76), respectively. Then, the optimal solution to problem (46) is computed at the CU via (44) and (47b). The key steps of the above decentralized DBP-based algorithm are summarized in Algorithm 2.
@@
1:$C$, $M$, NN, $K$, $P_{\max}$, $\alpha_{k}$, $\mathbf{\Sigma}_{k}$, $L_{k}^{\mathtt{Tx}}$, $L_{k}^{\mathtt{Rx}}$, $\theta_{ki}^{\mathtt{Tx}}$, $\phi_{ki}^{\mathtt{Tx}}$, $\theta_{kj}^{\mathtt{Rx}}$, $\phi_{kj}^{\mathtt{Rx}}$.
@@
5:   steps in Section V-C and store the results at the CU.
6:  Update each   $\mathbf{W}_{k}^{c}$   distributively   according to the
7:   steps in Section V-D and store it at the $c$-th DU.
8:  Update  $\mathbf{T}^{c}$  and   $\mathbf{R}_{k}$   distributively according to the
9:   steps in Section V-E,   and store them at the $c$-th DU
@@
[…]
Remark: Different from the decentralized algorithms in [51], which are mathematically equivalent to their centralized counterparts, the decentralized Algorithm 2 is not exactly equivalent to its centralized version, Algorithm 1. The inequivalence arises from the matrix inversion-free beamforming optimization introduced in Section V-D, as well as the approximation of the term $\hat{\mathbf{W}}$ in (74). Nevertheless, as will be demonstrated in Section VI, the centralized and decentralized algorithms achieve nearly identical performance.
V-F Complexity Analysis
For the centralized implementation given in Algorithm 1, the time complexity of a single BCA iteration is dominated by the updates of $\underline{\mathbf{W}}$ and $\mathbf{T}$, and is given by $\mathcal{O}(M^{3}T_{\rm bis}+M^{2}L_{k}^{\mathtt{Tx}}T_{\rm MM}^{\mathtt{Tx}})$. Here, $T_{\rm bis}$ denotes the number of iterations in the bisection search for updating $\underline{\mathbf{W}}$ described in Section IV-C, and $T_{\rm MM}^{\mathtt{Tx}}$ represents the average number of MM iterations required for updating $\mathbf{T}$. The complexity of the centralized implementation is proportional to $M^{3}$ and grows rapidly as the number of transmit FAs $M$ increases.
In comparison, the complexity of the decentralized implementation at the CU is given by $\mathcal{O}(K^{2}d^{3}+N^{3}K+T_{\rm MM}^{\mathtt{Rx}}(N^{2}L_{k}^{\mathtt{Rx}}+N(L_{k}^{\mathtt{Rx}})^{2}))$, and the complexity at each DU is given by $\mathcal{O}(NK^{2}d+K^{2}d^{2}+T_{\rm MM}^{\mathtt{Tx}}(M_{c}dL_{k}^{\mathtt{Tx}}+d(L_{k}^{\mathtt{Tx}})^{2}))$, where $T_{\rm MM}^{\mathtt{Tx}}$ and $T_{\rm MM}^{\mathtt{Rx}}$ represent the MM iterations for the update step of $\mathbf{T}$ and $\underline{\mathbf{R}}$, respectively. Notably, the complexities of the CU and each DU grow as a function of $M_{c}$, instead of $M$, and are significantly lower than those of the centralized implementation.
VI Simulation
In this section, we evaluate the performance of FA-assisted MU-MIMO networks optimized using the proposed centralized BCA-based algorithm in Algorithm 1 and its decentralized implementation in Algorithm 2. The centralized and decentralized algorithms are represented as “C” and “D”, respectively.
We denote the system with joint beamforming and transmit and receive FA position optimization as transmit and receive FA (TRFA). We compare the performance of TRFA with several baselines, specified as follows.
@@ VI Simulation
    FPA: The antenna arrays at the BS and users are fixed in position with a spacing of $\lambda/2$.
@@
    Transmit FA (TFA): The antennas at the users are fixed with a spacing of $\lambda/2$. The antennas at the BS are FAs.
@@
    Receive FA (RFA): The antennas at the BS are fixed with a spacing of $\lambda/2$, while the antennas at the users are FAs.
TABLE I: Key Simulation Parameters [34, 48, 51]
Parameter Value
Number of channel realizations $S=200$
Number of transmit FAs $M=64$
Number users $K=6$
User priority $\alpha_{k}=1$
Number of FAs at each user $N=4$
Number of parallel data streams $d=4$
Number of DUs $C=4$
Carrier freqency $f_{c}=28$ GHz
Carrier wavelength $\lambda=10.7$ $m$
Minimal distance between FAs $D=\lambda/2$
Noise power $\sigma_{k}^{2}=-90$ dBm
Transmit power budget $P_{\max}=30$ dBm
Miminum user distance from the BS $d_{\min}=100$ m
Maximum user distance from the BS $d_{\max}=300$ m
Distance from the BS to user $k$ $d_{k}^{2}\sim\mathcal{U}[d_{\min}^{2},d_{\max}^{2}]$
Pathloss exponent $\varrho=3.67$
Pathloss at reference distance $d_{0}=1$ m $T_{0}=-61.4$ dB
Elevation/Azimuth AoD $\theta_{kq}^{\mathtt{Tx}},\phi_{kq}^{\mathtt{Tx}}\sim\mathcal{U}[0,\pi)$
Elevation/Azimuth AoA $\theta_{kq}^{\mathtt{Rx}},\phi_{kq}^{\mathtt{Rx}}\sim\mathcal{U}[0,\pi)$
Number of transmit/receive paths $L_{k}^{\mathtt{Tx}}=L_{k}^{\mathtt{Rx}}=3$
Unless otherwise specified, the key simulation parameters follow the settings in Table I. The pathloss of user $k$ is calculated as $\kappa(d_{k})=T_{0}(d_{k}/d_{0})^{-\varrho}$, and the PRM is diagonal with entries following $\mathbf{\Sigma}_{k}]_{qq}\sim\mathcal{CN}(0,\kappa(d_{k})/L)$. Without the box-constrained movement mode, all transmit (and receive) FAs can move within a shared cuboid region, where the edge length is $\rho\lambda\sqrt{M}$ (and $\rho\lambda\sqrt{N}$), and the height is $2\rho\lambda$. With the box-constrained movement mode, the specific movable regions of each FA are detailed in Section IV-F.
VI-A Impact of Box-Constrained Movement Mode